# ABPT: Amended Backpropagation through Time with Partially Differentiable Rewards

### v1

@@ 4 Experiments
We address the following questions in this section:
  • How does ABPT improve performance on typical quadrotor tasks compared to baseline methods?
  • What distinctive advantages does ABPT exhibit in behavior?
  • What is the contribution of each individual component?
@@
4.1 Tasks for Evaluation
We conduct the evaluation on four quadrotor tasks, which involve different levels of complexity, where some of them are with non-differentiable rewards. A comprehensive descriptions of the observation and reward model are presented in Table 1.
@@ 4.1 Tasks for Evaluation
We use the quadrotor simulator VisFly Li et al. (2024) as our training environment, where the quadrotor dynamics are well implemented with automatic FOG computation achieved via Paszke et al. (2017).
@@
4.2 Comparison with Baseline Methods
We have compared our method with existing methods based on classic RL and differentiable physics. We present here the results of those on-policy methods, including PPO Schulman et al. (2017), BPTT, and SHAC Xu et al. (2022). As for off-policy methods, techniques such as SAC Haarnoja et al. (2018b) struggle to demonstrate stable training curves, leading us to suspect that training the value network off-policy poses significant challenges for our quadrotor tasks. Additionally, we implemented mixed gradient methods Suh et al. (2022); Gao et al. (2024) without the original source code, but these algorithms require extremely long training times due to the need for variance estimation at each step.
To ensure fair comparison, we implemented SHAC and BPTT by ourselves based on available source code, and adopt PPO from stable-baselines3 Raffin et al. (2021) in VisFly simulator. All algorithms used parallel differentiable simulations to accelerate training. We tuned all hyperparameters to achieve optimal performance, and kept the settings consistent across all experiments as possible as we can. All experiments were conducted on the same laptop with an RTX 4090 GPU and a 32-core 13th Gen Intel(R) Core(TM) i9-13900K processor, with 5 random seeds for validation of robustness. Given the different time-step metrics across the algorithms, we compare their performance in terms of wall-time as well. Further details of the settings can be found in our source code (to be released). Figure 5 provides reward curves of all methods during training.
[…]
4.3 Learning Rate Robustness
@@
4.4 Reward Robustness
5 Ablation Study
6 Discussion
@@
Gao et al. (2024) addresses the limitations of first-order gradients, such as bias under discontinuities and high variance in stiff contacts, by adaptive mixing policy gradients and first-order gradients at every time step. While probably effective in reducing the gradient bias from non-differentiable rewards, this method incurs significant computational overhead, limiting its practicality. In contrast, we mitigate the gradient bias by simply averaging the value gradient and the first-order gradient. This approach is far more efficient, requiring minimal overhead for evaluating the value gradients only at the initial and the last time steps.
@@
We also explored incorporating $k$-step value functions ($k=0,\ldots,N-1$) at each step within a finite horizon, following a similar approach to Suh et al. (2022). However, this led to significant fluctuations in the learning curves.
7 Conclusion
We present a novel approach to training policies for quadrotor tasks robustly. It effectively addresses several challenges associated with existing first-order gradient learning methods. Firstly, to address the non-differentiable rewards that cause bias in gradient backpropagation, we propose to combine the 0-step return with the N-step return in the objective function. This is equivalent to combining the value gradients with the first-order gradients for backpropagation, mitigating the bias issue. Secondly, we incorporate entropy regularization and re-sampling mechanisms to initialize visited states from the buffer to increase the randomization and encourage exploration during training.
We validated ABPT on four quadrotor tasks — hovering, tracking, landing, and racing -and compared them with existing learning algorithms. The results show that ABPT achieves faster and more stable training processes and converges higher rewards across all tasks. ABPT is also robust to the learning rate and different kinds of rewards. The ablation study also shows the effectiveness of each key component of our approach.

### v2

@@ 6 Experiments
We address the following questions in this section: 1) How does ABPT improve performance on typical quadrotor tasks compared to baseline methods? 2) What distinctive advantages does ABPT exhibit in behavior? 3) What is the contribution of each individual component?
6.1 Tasks for Evaluation
We conduct the evaluation on four quadrotor tasks, hovering, tracking, landing, racing, which involve different levels of complexity, where some of them are with non-differentiable rewards. Detailed task definitions are introduced in Appendix C.
6.2 Comparison with Baseline Methods
We have compared our method with existing methods based on classic RL and differentiable physics. We present here the results of those on-policy methods, including PPO [17], BPTT, and SHAC [35]. As for off-policy methods, techniques such as SAC [39] struggle to demonstrate stable training curves, leading us to suspect that training the value network off-policy poses significant challenges for our quadrotor tasks. Additionally, we implemented mixed gradient methods [11, 36] without the original source code, but these algorithms require extremely long training times due to the need for variance estimation at each step.
To ensure fair comparison, we implemented SHAC and BPTT by ourselves based on available source code, and adopt PPO from stable-baselines3 [40] in VisFly simulator. All algorithms used parallel differentiable simulations to accelerate training. We tuned all hyperparameters to achieve optimal performance, and kept the settings consistent across all experiments as possible as we can. All experiments were conducted on the same laptop with an RTX 4090 GPU and a 32-core 13th Gen Intel(R) Core(TM) i9-13900K processor, with 5 random seeds for validation of robustness. Given the different time-step metrics across the algorithms, we compare their performance in terms of wall-time as well. Further details of the settings can be found in our source code (to be released). Figure 4 provides reward curves of all methods during training.
[…]
@@ 6.2 Comparison with Baseline Methods
6.3 Additional results
We validate our trained policies in real-world experiments and provide a supplementary video demonstration. We further perform an ablation study on the key components of our approach ( Appendix D.1) and evaluate its robustness to different kinds of rewards ( Appendix D.2) as well as different learning rates ( Appendix D.3).
7 Conclusion
We present a novel approach ABPT to training policies for quadrotor tasks robustly. It effectively addresses the challenges from the partially non-differentiable rewards associated with existing first-order gradient learning methods. We validated ABPT on four quadrotor tasks — hovering, tracking, landing, and racing — and compared them with existing learning algorithms. The results show that ABPT achieves faster and more stable training processes and converges higher rewards across all tasks. ABPT is also robust to the learning rate and different kinds of rewards. The ablation study also shows the effectiveness of each key component of our approach. We also discuss the limitation of our approach in  Appendix B and present further discussion in  Appendix E.
Appendix A Proof
$\mathcal{G}_{\theta}^{t:t+N}\approx\mathcal{V}_{\theta|\phi}^{t}-(1-d)\gamma^{N}\mathcal{V}_{\theta|\phi}^{t+N+1}.$ (16)
Its value gradient is then given by
$\nabla_{\theta}^{[q]}\mathcal{G}_{\theta}^{t:t+N}=\nabla_{\theta}\mathcal{V}_{\theta|\phi}^{t}-(1-d)\gamma^{N}\nabla_{\theta}\mathcal{V}_{\theta|\phi}^{t+N+1}$ (17)
regardless of the differentiability of the rewards. Noting that, unlike [35], we specifically use action-value function $Q_{\phi}$ to compute the value to ensure $\mathcal{G}_{\theta}^{t:t+N}$ is differentiable with respect to $\theta$, which makes this derivative expression meaningful, otherwise the derivative would be zero if using $V_{\phi}$ solely with state input. Let $\nabla_{\theta}\mathcal{G}_{\theta}^{t:t+N}$ denote the first-order gradient of the accumulated reward. The average of the two gradients can be expressed as:
¯$\bar{\nabla}_{\theta}\mathcal{G}_{\theta}^{t:t+N}=\frac{1}{2}\left(\nabla^{[q]}_{\theta}\mathcal{G}_{\theta}^{t:t+N}+\nabla_{\theta}\mathcal{G}_{\theta}^{t:t+N}\right).$ (18)
Therefore, the difference between this gradient and the gradient (7) used in [35] is that the first-order gradients in (7) are combined with the value gradients. By leveraging this combination, our method remains effective in guiding the parameter updates toward the correct direction, when the first-order gradient is biased due to the non-differentiable rewards.
We conduct a simple experiment to assess the effectiveness of incorporating the 0-step return in addressing gradient bias. We deliberately detach parts of rewards in the hovering task (see Section 6.1) to mimic non-differentiable rewards,then back-propagate to compute gradient of network parameter. As shown in Figure 5, combining the 0-step return with the N-step return in the objective function (9) for training significantly reduces the model parameter residuals.
Appendix B Limitation
ABPT effectively enhances the efficiency and robustness of training processes utilizing analytical gradients, even in scenarios involving partially differentiable reward structures. However, while it significantly mitigates the gradient bias caused by non-differentiable reward components, it may still fail to fully eliminate extreme bias if the biased gradient is excessively large. In other words, if the non-differentiable rewards contribute too heavily to the overall reward, ABPT’s training performance may also degrade. Therefore, we strongly recommend that, when designing reward functions, priority should be given to incorporating smooth and differentiable variables to the greatest extent possible. In the following work, we will further explore how to adaptively mix the gradient while avoiding cost too much resources on mixture ratio computation.
Appendix C Task Defination
@@
We use the quadrotor simulator VisFly [29] as our training environment, where the quadrotor dynamics are well implemented with automatic FOG computation achieved via [32].A comprehensive description of observation and reward structure is presented in Table 1.
Table 1: Observations and rewards used in benchmark quadrotor tasks
Environments Observation Reward Function
Hovering state & $\hat{p}$ $c-k_{1}\left\|p-\hat{p}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|$ (fully DIFF)
Tracking state & next 10 $\hat{p}_{i=1\sim 10}$ $c-k_{1}\left\|p-\hat{p}_{0}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|$ (fully DIFF)
Landing state & $\hat{p}$ $-k_{1}f^{+}\big{(}\left\|p_{xy}-\hat{p}_{xy}\right\|\big{)}+k_{2}f^{+}\big{(}\left\|v_{z}-\hat{v}_{z}\right\|\big{)}+k_{3}s$ (partially DIFF)
Racing state & next 2 $\hat{p}_{i=1,2}$ of gates $c-k_{1}\left\|p-\hat{p}_{0}\right\|-k_{2}\left\|q-\hat{q}\right\|-k_{3}\left\|v\right\|-k_{4}\left\|\omega\right\|+k_{5}s$ (partially DIFF)
  • 1
    $c$ represents a small constant used to ensure the agent remains alive. $k_{i}$ denotes constant weights for different reward contributions, with these weights being distinctly defined for each task. $s$ is a boolean variable that indicates whether the task is successfully completed. The state comprises position ($p$), orientation ($q$), linear velocity ($v$), and angular velocity ($\omega$). $f^{+}(\cdot)$ and $\hat{(\cdot)}$ are separately an increasing mapping function used to normalize the reward and target status. DIFF is abbreviation for differentiable. All the action types are individual thrusts.
Appendix D Additional Experiment
D.1 Ablation Study
D.2 Reward Robustness
D.3 Learning Rate Robustness
@@
Appendix E Discussion
@@
The authors in [36] address the limitations of first-order gradients, such as bias under discontinuities and high variance in stiff contacts, by adaptive mixing policy gradients and first-order gradients at every time step. While probably effective in reducing the gradient bias from non-differentiable rewards, this method incurs significant computational overhead, limiting its practicality. In contrast, we mitigate the gradient bias by simply averaging the value gradient and the first-order gradient. This approach is far more efficient, requiring minimal overhead for evaluating the value gradients only at the initial and the last time steps.
@@
We also explored incorporating $k$-step value functions ($k=0,\ldots,N-1$) at each step within a finite horizon, following a similar approach to [11]. However, this led to significant fluctuations in the learning curves.
Appendix F Training Hyperparameters
Tables 2$\sim$5 contain the parameters for the baseline experiments, Tables 6$\sim$8 for the ablation experiments, and Table 9 for the reward robustness experiments.
Table 2: Hyerparameters of SHAC
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.01 0.002
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 10
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 0.005 0.005
decay learning rate False False False True
value estimation factor $\lambda$ 0.95 0.95 0.95 0.95
horizon length $H$ 96 96 96 96
Optimizer Adam Adam Adam Adam
Table 3: Hyerparameters of PPO
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.001 0.0002 0.0005 0.001
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
minibatch size 25600 25600 25600 51200
training critic steps per minibatch 5 5 5 5
weight decay 0.00001 0.00001 0.00001 0.00001
GAE $\lambda$ 1 1 1 1
Optimizer Adam Adam Adam Adam
Table 4: Hyerparameters of BPTT
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.005 0.002
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
weight decay 0.00001 0.00001 0.00001 0.00001
decay learning rate False False False True
horizon length $H$ 256 256 256 512
Optimizer Adam Adam Adam Adam
Table 5: Hyerparameters of ABPT
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.01 0.01
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 10
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 0.005 0.005
decay learning rate False True False True
value estimation factor $\lambda$ 0.95 0.95 0.95 0.95
horizon length $H$ 96 96 96 96
replay buffer size 1000000 1000000 1000000 50000
Optimizer Adam Adam Adam Adam
Table 6: Hyerparameters of ABPT no 0-step Value
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.01 0.01
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 10
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 0.005 0.005
decay learning rate False True False True
value estimation factor $\lambda$ 0.95 0.95 0.95 0.95
horizon length $H$ 96 96 96 96
replay buffer size 1000000 1000000 1000000 50000
Optimizer Adam Adam Adam Adam
Table 7: Hyerparameters of ABPT no Entropy
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.002 0.002
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 10
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 0.005 0.005
decay learning rate False True False True
value estimation factor $\lambda$ 0.95 0.95 0.95 0.95
horizon length $H$ 96 96 96 96
replay buffer size 1000000 1000000 1000000 50000
Optimizer Adam Adam Adam Adam
Table 8: Hyerparameters of ABPT no 0-step Value no Entropy
Hovering Tracking Landing Racing
learning rate $\alpha$ 0.01 0.01 0.002 0.002
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 10
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 0.005 0.005
decay learning rate False True False True
value estimation factor $\lambda$ 0.95 0.95 0.95 0.95
horizon length $H$ 96 96 96 96
replay buffer size 1000000 1000000 1000000 50000
Optimizer Adam Adam Adam Adam
Table 9: Hyerparameters of ABPT upon Velocity-based Reward in Racing
ABPT SHAC BPTT PPO
learning rate $\alpha$ 0.02 0.02 0.002 0.0002
number of parallel environments $n$ 100 100 100 100
discount factor $\gamma$ 0.99 0.99 0.99 0.99
training critic steps per minibatch 10 10 10 5
weight decay 0.00001 0.00001 0.00001 0.00001
target critic update factor $\tau$ 0.005 0.005 - -
decay learning rate True True True -
value estimation factor $\lambda$ 0.95 0.95 - -
horizon length $H$ 96 96 512 -
replay buffer size 50000 - - -
minibatch size - - - 51200
GAE - - - 1
Optimizer Adam Adam Adam Adam
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.